# Emotion-Coherent Reasoning for Multimodal LLMs via Emotional Rationale Verifier

### v1

@@ 3.3 Explanation Reward
EMER-SFT     33.06 29.67 70.66 38.95 25.21     31.47 24.28 56.49 38.93 30.86
EMER-SFT (7B)     44.26 38.58 67.38 50.22 33.48     40.68 27.58 42.27 48.35 50.13
Emotion-LLaMA (7B)     - - - - -     31.70 21.39 42.14 48.71 38.54
@@
R1-Omni (Zhao, Wei, and Bo 2025)     37.87 31.80 48.09 57.68 40.04     33.49 28.09 41.58 65.44 56.94
Baseline     32.41 27.27 40.98 56.83 38.66     32.37 27.57 40.16 66.17 56.10
Baseline (7B)     44.97 40.22 55.19 65.36 47.27     43.34 38.54 46.94 77.09 73.73
@@ 4.3 Implementation Details
We employ HumanOmni 0.5B model (Zhao et al. 2025b) as a backbone model. Our ERV is based on the RoBERTA-100M trained on the GoEmotions Dataset. Our training pipeline consists of three components: SFT, ERV module, and GRPO training. The main model is first fine-tuned on the EMER dataset for 5 epochs with a learning rate of $2\mathrm{e}^{-5}$, a cosine scheduler, a warmup ratio of 0.03, and a batch size of 32. The ERV module is trained separately for 10 epochs with a batch size of 64 and a learning rate of $3\mathrm{e}^{-5}$. Those are trained on 8 NVIDIA RTX 3090 GPUs. For GRPO training, we use a batch size of 16 with a gradient accumulation step of 2, generating $G=16$ samples per input, while applying a KL divergence penalty with coefficient $\beta=0.04$ and fixing the sampling temperature at 0.3 during inference. This stage is conducted on 8 NVIDIA A100 GPUs. Appendix E shows more training details.
@@ 5.1 Main Results
As shown in Table 3.3, both EMER-SFT and R1-Omni exhibit low FCR, achieving only 29.67% and 31.80%, respectively. This indicates that although R1-Omni significantly improves emotion recognition accuracy on WAR, from 38.95% to 57.68% compared to EMER-SFT, 44.87% of its predictions (almost half) are not supported by emotionally aligned explanations. Furthermore, its ability to independently produce explanations aligned with the target emotions improves only marginally from 33.06% to 37.87%. In contrast, our model trained with an Explanation reward not only achieves comparable recognition performance to R1-Omni but also demonstrates substantial improvements in both FCR and EEA. FCR increases from 29.67% to 43.57% and also EEA increases from 33.06% to 48.93%. A similar tendency is observed on the DFEW, where FCR improves from 24.28% to 40.21% and EEA improves from 31.47% to 45.82%.
We also compared our model with EMER-SFT (7B) and Emotion-LLaMA (7B), both of which use 7B-sized backbones but have different architectures. In the evaluation on the DFEW dataset, both models exhibited low performance, achieving EEA scores of 31.47% and 31.70%, and FCR scores of 24.28% and 21.39%, respectively. These results indicate that SFT training on existing multimodal video–reasoning pair datasets yields unreliable emotion reasoning performance regardless of the model architecture. In contrast, despite using a smaller 0.5B LLM size, our model outperforms both baselines in terms of FCR and EEA, while also achieving noticeably better emotion recognition accuracy. When applied to 7B-scale models, our method shows comparable recognition accuracy to Emotion-LLaMA fine-tuned solely for emotion label prediction, while achieving 45.95% FCR and 50.88% EEA.
About EPC, EMER-SFT shows significant performance compared to RL-based trained models. Namely, the final predicted answer is derived based on the generated explanations. However, when trained with GRPO on answer and format reward, this consistency deteriorates, as the model tends to focus solely on producing the final answer, leading to a degradation in EPC. In contrast, our proposed Explanation Reward enables the model to maintain this consistency, achieving comparable EPC performance to EMER-SFT (70.66% vs. 68.58% in MAFW and 56.49% vs. 58.63% in DFEW). This indicates that our Explanation Reward effectively guides the model to generate explanations and predictions in a coherent and aligned manner.
@@ 6 Conclusion
We proposed training an MLLM for the MER task utilizing our proposed ERV and Explanation Reward, aiming to enhance the coherence between emotion recognition and explanation. Compared to prior approaches, where coherence between predictions and explanations among correctly predicted samples remained around half (55.13% on MAFW and 42.92% on DFEW), our method significantly improves this alignment, achieving 75.79% and 59.77%.Furthermore, human evaluations confirmed that our approach produces more emotionally and contextually grounded emotion explanations for video data. As a result, our work enables emotionally coherent and trustworthy explanations alongside accurate emotion recognition, representing a significant advancement toward emotionally intelligent human–computer interaction.

### v2

@@ 3.3 Explanation Reward
EMER-SFT     33.06 29.67 70.66 38.95 25.21     35.14 32.63 64.30 38.93 30.86
EMER-SFT (7B)     44.26 38.58 67.38 50.22 33.48     44.86 42.21 57.41 48.33 39.16
Emotion-LLaMA (7B) (Cheng et al. 2024a)     - - - - -     41.15 32.56 57.69 48.71 38.54
@@
R1-Omni (Zhao, Wei, and Bo 2025)     37.87 31.80 48.09 57.68 40.04     43.69 37.99 51.91 65.83 56.27
Baseline     32.41 27.27 40.98 56.83 38.66     41.66 36.39 48.67 66.17 55.33
Baseline (7B)     44.97 40.22 55.19 65.36 47.27     55.39 52.01 60.70 76.67 69.69
@@ 4.3 Implementation Details
We employ HumanOmni 0.5B model (Zhao et al. 2025b) as a backbone model. Our ERV is based on the RoBERTA-100M trained on the GoEmotions Dataset. Our training pipeline consists of three components: SFT, ERV module, and GRPO training. The main model is first fine-tuned on the EMER dataset for 5 epochs with a learning rate of $2\mathrm{e}^{-5}$, a cosine scheduler, a warmup ratio of 0.03, and a batch size of 32. The ERV module is trained separately for 10 epochs with a batch size of 64 and a learning rate of $3\mathrm{e}^{-5}$. Those are trained on 8 NVIDIA RTX 3090 GPUs. For GRPO training, we use a batch size of 16 with a gradient accumulation step of 2, generating $G=16$ samples per input, while applying a KL divergence penalty with coefficient $\beta=0.04$ and fixing the sampling temperature at 0.3 during inference. For the 7B model, we trained with $G=4$ instead. This stage is conducted on 8 NVIDIA A100 GPUs. Appendix E shows more training details.
@@ 5.1 Main Results
As shown in Table 3.3, both EMER-SFT and R1-Omni exhibit low FCR, achieving only 29.67% and 31.80%, respectively. This indicates that although R1-Omni significantly improves emotion recognition accuracy on WAR, from 38.95% to 57.68% compared to EMER-SFT, 44.87% of its predictions (almost half) are not supported by emotionally aligned explanations. Furthermore, its ability to independently produce explanations aligned with the target emotions (EEA) improves only marginally from 33.06% to 37.87%. In contrast, our model trained with an Explanation reward not only achieves comparable recognition performance to R1-Omni but also demonstrates substantial improvements in both FCR and EEA. FCR increases from 29.67% to 43.57% and also EEA increases from 33.06% to 48.93%. A similar tendency is observed on the DFEW, where FCR improves from 32.63% to 53.75% and EEA improves from 35.14% to 59.45%.
We also compared our model with EMER-SFT (7B) and Emotion-LLaMA (7B), both of which use 7B-sized backbones but have different architectures. In the evaluation on the DFEW dataset, both models exhibited low performance, achieving EEA scores of 44.86% and 41.15%, and FCR scores of 42.21% and 32.56%, respectively. These results indicate that SFT training on existing multimodal video–reasoning pair datasets yields unreliable emotion reasoning performance regardless of the model architecture. In contrast, despite using a smaller 0.5B LLM size, our model outperforms both baselines in terms of FCR and EEA, while also achieving noticeably better emotion recognition accuracy. When applied to 7B-scale models, our method shows comparable recognition accuracy to Emotion-LLaMA fine-tuned solely for emotion label prediction, while achieving 62.06% FCR and 65.50% EEA.
About EPC, EMER-SFT shows significant performance compared to RL-based trained models. Namely, the final predicted answer is derived based on the generated explanations. However, when trained with GRPO on answer and format reward, this consistency deteriorates, as the model tends to focus solely on producing the final answer, leading to a degradation in EPC. In contrast, our proposed Explanation Reward enables the model to maintain this consistency, achieving performance that is comparable to or even surpasses that of EMER-SFT (70.66% vs. 68.58% in MAFW and 64.30% vs. 73.06% in DFEW). This indicates that our Explanation Reward effectively guides the model to generate explanations and predictions in a coherent and aligned manner.
@@ 6 Conclusion
We proposed training an MLLM for the MER task utilizing our proposed ERV and Explanation Reward, aiming to enhance the coherence between emotion recognition and explanation. Compared to prior approaches, where coherence between predictions and explanations among correctly predicted samples remained around half (55.13% on MAFW and 42.92% on DFEW), our method significantly improves this alignment, achieving 75.79% and 79.87%. Furthermore, human evaluations confirmed that our approach produces more emotionally and contextually grounded emotion explanations for video data. As a result, our work enables emotionally coherent and trustworthy explanations alongside accurate emotion recognition, representing a significant advancement toward emotionally intelligent human–computer interaction.